# Cognitive Memory in Large Language Models

### v1

@@ 4.3.2 Memory Management Integrated with the OS System
Prabhu et al. (2024) introduces a dynamic memory management system called vAttention for LLM (Language Model) service systems. vAttention leverages the operating system’s support for virtual memory and demand paging to reduce the programming burden and performance loss associated with traditional PagedAttention methods.
@@
Lin et al. (2024) introduces Infinite-LLM, a novel LLM (Large Language Model) serving system designed to address the highly dynamic context length management in LLM requests. The paper first analyzes the computational characteristics of LLM models, pointing out the limitations of traditional static model parallelization and KVCache scheduling methods when dealing with dynamic context lengths. To tackle these issues, the paper proposes three innovations: the DistAttention mechanism, which distributes attention computation and KVCache across the entire GPU cluster to improve cluster throughput; the liability mechanism, which allows borrowing memory from other instances to handle large context tasks and increase generation throughput; and gManager, which is used for global planning and coordination of request and KVCache allocation.
@@ 4.3.3 Shared Attention
Large language models typically precompute KV caches to speed up prefilling when processing multi-text-chunk inputs, but Yao et al. (2024a) is only effective when the text chunk is a prefix of the input. When the text chunk is not a prefix, the precomputed KV cache cannot be directly used due to the lack of cross-attention with preceding text, limiting cache reuse. To address this issue, the authors propose a scheme called CacheBlend. CacheBlend reuses precomputed KV caches regardless of whether the text chunk is a prefix or not. It selectively recomputes the KV values of a small subset of tokens to partially update the reused KV caches and supplement the cross-attention. The delay for recomputing these tokens can be pipelined with the retrieval of KV caches, allowing CacheBlend to store caches in slower, higher-capacity devices without increasing inference delay. This approach maintains the same generation quality as full prefill while improving efficiency.
@@
4.3.4 The Relationship between KV Cache Compression and Attention.
@@
The attention mechanism is a core component of the Transformer architecture, and it has enabled the remarkable success of large-scale language models. However, the theoretical underpinnings of the attention mechanism are not yet fully understood, especially its non-convex optimization dynamics. In Ataee Tarzanagh et al. (2023), the authors investigate the pioneering softmax attention model $f(X)=\langle Xv,\text{softmax}(XWp)\rangle$, where $X$ is a sequence of tokens, and $(v,W,p)$ are trainable parameters. They prove that running gradient descent on $p$, or equivalently on $W$, leads to directional convergence towards a maximum-margin solution that can distinguish between locally optimal tokens and non-optimal tokens. This explicitly formalizes attention as an optimal token selection mechanism. Notably, the authors’ results apply to general data and precisely characterize the optimality of tokens in terms of the value embeddings $Xv$ and the geometry of the problem. The authors also provide a broader analysis of the regularization path, establishing the maximum-margin property of attention even in the presence of a nonlinear prediction head. When $v$ and $p$ are optimized simultaneously and a logistic loss is used, the authors determine the conditions under which the regularization path directionally converges to their respective hard-margin support vector machine solutions, where $v$ separates the input features based on the labels.
5 Parameters-Based Memory
Both text-based memory and KV cache-based memory are merely storage of historical information, and they do not become the model’s true "memory." In this paper, we introduce a method of memory parameterization, which transforms previous memories into parameters as part of the model itself.
We categorize the strategies of memory parameterization into three types: LoRA, Test Time Training (TTT), and Mixture of Experts (MoE), and introduce each of them respectively.
@@ 6 Hidden-State-Based Memory
In addition to retaining text memory, KV cache, and parametrized memory, some researchers have combined the idea of hidden states in RNNs with current methods. These methods can be categorized into three types: methods based on the chunk mechanism, recurrent transformer methods, and methods based on Mamba.
@@ 6.1 Chunk Mechanism
The previous process leads to a fixed heuristic pattern of discarding tokens, which is not ideal as the model may naively retain less valuable tokens while discarding important ones. To solve this, the authors dynamically select important tokens to retain by tracking their average attention scores using exponential moving average (EMA). They discard tokens with lower EMA scores. Specifically, given a hyperparameter $\gamma\in[0,1]$ and attention score vector $\mathbf{s}_{k}^{(t)}$ at time step $t$, the average attention score $\mu^{(t)}$ is updated as:
@@
Bertsch et al. (2023) introduces a method called Unlimiformer, which can accept input texts of arbitrary length during testing. Unlimiformer constructs the hidden states of the input text using a k - nearest neighbors index. It then queries the k - nearest neighbors index in the standard cross - attention heads of each decoder layer, using the k - nearest neighbor distances as attention scores and focusing only on the top - k input tokens. In this way, Unlimiformer is able to extend existing encoder - decoder Transformer models to accept inputs of arbitrary length. This method not only has low computational and memory overhead, but also accurately approximates global attention.
@@
An et al. (2024) introduces a novel training-free framework called Dual Chunk Attention (DCA) for extending the context window of large language models (LLMs). The paper first describes the three components of DCA: intra-chunk attention, inter-chunk attention, and successive chunk attention. These attention mechanisms help the model effectively capture both long-range and short-range dependencies when processing long sequences.
@@ 6.2 Recurrent Transformer
The Recurrent Transformer is a hybrid model that combines the strengths of Recurrent Neural Networks (RNNs), such as their ability to retain information when processing long sequences, with the parallel processing and global dependency capturing capabilities of the Transformer model. This type of model typically includes an embedding layer that maps input words to high-dimensional vectors, a recurrent layer such as LSTM or GRU to capture long-term dependencies, and self-attention mechanisms to process all elements of the sequence in parallel. Additionally, feed-forward networks perform nonlinear transformations on the outputs of the attention layers, while layer normalization helps stabilize the training process. Residual connections are formed between the outputs and inputs of each sublayer to avoid the vanishing gradient problem. The Recurrent Transformer may include encoder and decoder components, and some models adopt hierarchical structures to handle sequences at different levels. In the model, information flows through the recurrent layer to capture long-term dependencies, then through the attention mechanism to understand the global relationships between elements, and finally through the feed-forward network for further transformation. Different variants of the Recurrent Transformer may adjust these components or introduce new mechanisms according to specific applications and research objectives to enhance performance.

### v2

@@ 4.3.2 Memory Management with OS System
Prabhu et al. (2024) introduces a dynamic memory management system called vAttention for LLM (Language Model) service systems. attention leverages the operating system’s support for virtual memory and demand paging to reduce the programming burden and performance loss associated with traditional PagedAttention methods.
@@
Lin et al. (2024) introduces Infinite-LLM, a novel LLM (Large Language Model) serving system designed to address the highly dynamic context length management in LLM requests. The paper first analyzes the computational characteristics of LLM models, pointing out the limitations of traditional static model parallelization and KVCache scheduling methods when dealing with dynamic context lengths. To tackle these issues, the paper proposes three innovations: the DistAttention mechanism, which distributes attention computation and KVCache across the entire GPU cluster to improve cluster throughput; the liability mechanism, which allows borrowing memory from other instances to handle large context tasks and increases generation throughput; and gManager, which is used for global planning and coordination of request and KVCache allocation.
@@ 4.3.3 Shared Attention
Large language models typically precompute KV caches to speed up prefilling when processing multi-text-chunk inputs, but Yao et al. (2024a) is only effective when the text chunk is a prefix of the input. When the text chunk is not a prefix, the precomputed KV cache cannot be directly used due to the lack of cross-attention with the preceding text, limiting cache reuse. To address this issue, the authors propose a scheme called CacheBlend. CacheBlend reuses precomputed KV caches regardless of whether the text chunk is a prefix or not. It selectively recomputes the KV values of a small subset of tokens to partially update the reused KV caches and supplement the cross-attention. The delay for recomputing these tokens can be pipelined with the retrieval of KV caches, allowing CacheBlend to store caches in slower, higher-capacity devices without increasing inference delay. This approach maintains the same generation quality as full prefill while improving efficiency.
@@
4.3.4 Relationship between KV Cache Compression and Attention
The attention mechanism is a core component of the Transformer architecture and has enabled the remarkable success of large-scale language models. However, the theoretical underpinnings of the attention mechanism are not yet fully understood, especially its non-convex optimization dynamics. In Ataee Tarzanagh et al. (2023), the authors investigate the pioneering softmax attention model $f(X)=\langle Xv,\text{softmax}(XWp)\rangle$, where $X$ is a sequence of tokens and $(v,W,p)$ are training parameters. They prove that running gradient descent on $p$, or equivalently on $W$, leads to directional convergence towards a maximum-margin solution that can distinguish between locally optimal tokens and non-optimal tokens. This explicitly formalizes attention as an optimal token selection mechanism. In particular, the authors’ results apply to general data and precisely characterize the optimality of tokens in terms of the value embeddings $Xv$ and the geometry of the problem. The authors also provide a broader analysis of the regularization path, establishing the maximum-margin property of attention even in the presence of a non-linear prediction head. When $v$ and $p$ are optimized simultaneously and logistic loss is used, the authors determine the conditions under which the regularization path converges in the direction of their respective hard-margin support vector machine solutions, where $v$ separates the input features based on the labels.
5 Parameters-Based Memory
@@
Both text-based memory and KV cache-based memory are merely storage of historical information, and they do not become the model’s true "memory." In this section, we introduce a method of memory parameterization, which transforms previous memories into parameters as part of the model itself.
We categorize the strategies of memory parameterization into three types: LoRA, Test Time Training (TTT), and Mixture of Experts (MoE), and introduce each of them respectively, as shown in Figure 8.
@@ 6 Hidden-State-Based Memory
In addition to retaining text memory, KV cache, and parametrized memory, some researchers have combined the idea of hidden states in RNNs with current methods. These methods can be categorized into three types: methods based on the chunk mechanism, recurrent transformer methods, and methods based on Mamba. Figure 9 shows the overall structure.
@@ 6.1 Chunk Mechanism
The previous process leads to a fixed heuristic pattern of discarding tokens, which is not ideal as the model may naively retain less valuable tokens while discarding important ones. To solve this, the authors dynamically select important tokens to retain by tracking their average attention scores using exponential moving averages (EMA). They discard tokens with lower EMA scores. Specifically, given a hyperparameter $\gamma\in[0,1]$ and attention score vector $\mathbf{s}_{k}^{(t)}$ at time step $t$, the average attention score $\mu^{(t)}$ is updated as:
@@
Bertsch et al. (2023) introduces a method called Unlimiformer, which can accept input texts of arbitrary length during testing. Unlimiformer constructs the hidden states of the input text using a k - k-nearest neighbors index. It then queries the k - nearest neighbors index in the standard cross-attention heads of each decoder layer, using the k - nearest neighbor distances as attention scores and focusing only on the top - k input tokens. In this way, Unlimiformer is able to extend existing encoder-decoder Transformer models to accept inputs of arbitrary length. This method not only has low computational and memory overhead but also accurately approximates global attention.
@@
An et al. (2024) introduces a novel training-free framework called Dual Chunk Attention (DCA) for extending the context window of large language models (LLMs). The paper first describes the three components of DCA: intra-chunk attention, inter-chunk attention, and successive-chunk attention. These attention mechanisms help the model effectively capture both long-range and short-range dependencies when processing long sequences.
@@ 6.2 Recurrent Transformer
The Recurrent Transformer is a hybrid model that combines the strengths of Recurrent Neural Networks (RNNs), such as their ability to retain information when processing long sequences, with the parallel processing and global dependency-capturing capabilities of the Transformer model. This type of model typically includes an embedding layer that maps input words to high-dimensional vectors, a recurrent layer such as LSTM or GRU to capture long-term dependencies, and self-attention mechanisms to process all elements of the sequence in parallel. Additionally, feed-forward networks perform nonlinear transformations on the outputs of the attention layers, while layer normalization helps stabilize the training process. Residual connections are formed between the outputs and inputs of each sublayer to avoid the vanishing gradient problem. The Recurrent Transformer may include encoder and decoder components, and some models adopt hierarchical structures to handle sequences at different levels. In the model, information flows through the recurrent layer to capture long-term dependencies, then through the attention mechanism to understand the global relationships between elements, and finally through the feed-forward network for further transformation. Different variants of the Recurrent Transformer may adjust these components or introduce new mechanisms according to specific applications and research objectives to enhance performance.
@@ 7 Discussion
In the current classification of memory mechanisms, text-based memory and memory rooted in model parameters are categorized as long-term memory. These forms of memory can endure across multiple session cycles. In contrast, memory based on KV caches and memory based on hidden states are classified as short-term memory. They exist solely within a single session cycle and vanish once the session concludes.
When compared to human memory mechanisms, current Large Language Models (LLMs) still exhibit significant deficiencies in the realm of information retrieval. A key distinction lies in the fact that humans possess an exceptional ability to summarize and generalize information. Humans can distill complex and intricate matters into simple and clear principles. This process not only significantly reduces the “space” required to store information but also greatly enhances the applicability of memory. The generalized principles can be readily activated and applied across a wide array of different situations, demonstrating remarkable flexibility and generalization capabilities. Another notable difference is that LLMs lack an inherent ability to forget. Whether it is memory based on text, KV caches, or hidden states, the models will retain all previous information intact in the absence of active human intervention. This leads to the continuous accumulation and expansion of information, which may negatively impact the performance and efficiency of the models. For instance, it can increase storage burdens and reduce retrieval speeds. Humans also possess the ability to integrate and assimilate knowledge, which represents an advanced form of generalization. When new memories are introduced that are related to or in conflict with existing memories, humans can skillfully integrate the new and old memories to form new memories that replace the original ones. This ability allows human memory to dynamically adapt to new knowledge and information, maintaining its accuracy and relevance.
Among the current methods, short-term memory based on KV caches and hidden states is inherently limited, and Their storage and retrieval capabilities are confined to the current session cycle. On the other hand, text-based memory has theoretical upper limits in terms of both storage capacity and retrieval efficiency. These limitations make it challenging to meet the growing demands of information processing. Therefore, memory parameterization has emerged as a promising direction for the future development of LLM cognitive memory, as it closely resembles human memory mechanisms. The goal of memory parameterization is to enable models to acquire abilities such as summarizing information, integrating knowledge, and selectively forgetting details, akin to human cognitive processes, thereby improving the efficiency of information processing and storage. Significant progress has been made in the field of memory parameterization. Methods such as LoRA, MOE, TTT, and the design of new modules have been explored. Future research should focus on effectively transforming memory into model parameters without compromising the original memory and without significantly increasing the number of model parameters. This challenging yet highly promising area of research holds the potential to drive significant advancements in LLM cognitive memory, ultimately bringing these models closer to human cognitive capabilities and memory functions.